The weights of bags of baby carrots are normally​ distributed, with a mean of 28 ounces and a standard deviation of 0.33 ounce. Bags in the upper​ 4.5% are too heavy and must be repackaged. What is the most a bag of baby carrots can weigh and not need to be​ repackaged?
Answer #1

Given that,

mean = = 28

standard deviation = = 0.33

Using standard normal table,

P(Z > z) =4.5 %

= 1 - P(Z < z) = 0. 045

= P(Z < z ) = 1 - 0.045

= P(Z < z ) = 0.955

= P(Z < 1.70 ) = 0.  

z = 1.70

Using z-score formula  

x = z +

x =1.70 *0.33+28

x = 28.561

x = 29